OpenStudy (anonymous):

@Reaper534 A football team lost 9 yards on first down and 4 yards on second down. It gained 5 yards on third down. How many yards does it need to gain on fourth down to have a 10-yard gain from its starting position? A. 2 yd B. 8 yd C. 13 yd D. 18 yd

OpenStudy (texaschic101):

x -9 - 4 + 5 = 10 x - 13 + 5 = 10 x - 8 = 10 x = 10 + 8 x = 18 not sure if I set it up right...you might need to get a second opinion

OpenStudy (texaschic101):

because when I keep looking at it...if he lost 13 yards, then gained 5...that would put him at 8...needing only 2 yds. Not sure about this....

OpenStudy (texaschic101):

@Reaper534 ...any thoughts.....

OpenStudy (anonymous):

@Reaper534 have anything to say

OpenStudy (texaschic101):

I am gonna go with 2 yds

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@texaschic101 are u sure

OpenStudy (anonymous):

no seems good

OpenStudy (anonymous):

yeah 2 yards

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

yeah 2 yards it is

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

@mathmale @mathstudent55

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

no i think its D because it said they lost 9 and 4 yard leaves -13 and gaind 5 mean they have -8 yards

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

and it asked how many yards are needed to to get 10 yards

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

hotchellerae i think it is D

OpenStudy (anonymous):

did you see what i put up there

OpenStudy (anonymous):

D

OpenStudy (anonymous):